A monopolist has cost function TC=10+2Q, where TC is the total cost of producing Q units of output. Demand in this market is given by the equation Q=14-P , where P stands for the price. Calculate the profite that the monopolist will be marking.

1
Expert's answer
2021-12-21T11:55:48-0500

TC= 10+2Q

MC= 2

Q= 14-P

P= 14-Q

TR= P*Q

= Q(14-Q)

=14Q-Q2

MR= 14-2Q

"\therefore" Profit Maximization output = MR=MC

14-2Q=2

2Q=12

Q=6

"\Pi" =TR-TC

=14(6)-62

=48

TC= 10+2(6)

= 22

"\therefore" "\Pi" =48-22

= 26